OpenStudy (anonymous):

Jim is 8 years old, and his Uncle Bill is 5 1/2 times older than he his. What is his Uncle Bill's age?

OpenStudy (igreen):

First find out what 5 * 8 is. 5 * 8 = 40. Then find out what half of 8 is. 1/2 * 8 = 4. Now add: 40 + 4 = 44. So Bill's Uncle Bill is 44 years old.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Jammer123